I pick up 3 cards from a standard deck, one after another, keeping each card that I pick (without replacement) What is the probability that as least 2 of the 3 cards are the same suit?
B) There are 10 cards in a hat, 4 red and the rest black. You pull out 2 cards without looking. What is the probability that they are the same color? Would this be (Prob of red)= 4/10*3/9+(Prob of all black) 6/10*5/9=42/90= 46%

Answer by stanbon(75887)

I pick up 3 cards from a standard deck, one after another, keeping each card that I pick (without replacement) What is the probability that as least 2 of the 3 cards are the same suit?
Prob(2 or 3 are the same suit) = 1 - P(none are the same suit)
Prob(none) = ?
1st: pick any card (52/52)
2nd: pick a card not of that suit (39/51)
3rd: pick a card of neither picked suit (26/50)
Therefore ? = 1 - (39/51)(26/50) = 1-0.39765... = 0.6024..
=======================


B) There are 10 cards in a hat, 4 red and the rest black. You pull out 2 cards without looking. What is the probability that they are the same color? Would this be (Prob of red)= 4/10*3/9+(Prob of all black) 6/10*5/9=42/90= 46%
------
P(2 red or 2 black) = [4C2+6C2]/10C2 = [6 + 15]/45 = 21/45
============
Cheers,
Stan H.
------------------